You are given the following information for Calvani Pizza Co.: sales = $41200; costs = $21601; addition to retained earnings = $7046; dividends paid = $3793; interest expense = $2391; tax rate = 30 percent. Calculate the depreciation expense.
Answer #1

Net income = Retained earnings + Dividend

Net income = 7,046 + 3,793 = 10,839

Tax expense = 30%

Since, net income calculated after deducting tax expense form pre tax income, hence

Pre tax income = 10,839 * 70 / 100 = 15,484.29

Now

Pre tax income = Sales - Cost - Depreciation - interest expense

Depreciation = Sales - Cost - interest expense - pre tax income

Depreciation = 41,200 - 21,601 - 2,391 - 15,484.29 = 1,723.71